OpenStudy (anonymous):

If you double a number then add 28 you get 4/9 of the original number. What is the original number?

OpenStudy (anonymous):

let's x be your number, then: 2x+28=4/9 x solve it for x, please

OpenStudy (anonymous):

I get -108/7

OpenStudy (anonymous):

And I know it's wrong.

OpenStudy (anonymous):

2x-4/9x=-28 (18-4)/9 x=-28 14x=-28 * 8 x=-252/14 x=-18

OpenStudy (anonymous):

do you see when you made a mistake? fractions....?

OpenStudy (anonymous):

Fractions confuse the hell out of me.

OpenStudy (anonymous):

just remember that when you + or "-" you need common denominator... you do not need to worry about it when you multiply or divide. keep doing exercises - it's helps to get use to it :) Good Luck!